# A Decade of News Forum Interactions: Threaded Conversations, Signed Votes, and Topical Tags

### v1

@@ 3.3 Internal Consistency
As noted in the main text, we provide a separate file listing all comment identifiers associated with potential data inconsistencies. These include: (1) comments referencing article identifiers that are no longer available in the articles metadata (nor accessible on the DerStandard website), and (2) comments with timestamps preceding those of their associated articles. While these cases represent a small fraction of the data, they may affect specific analyses, such as estimating user lifespans. We recommend that users assess whether these entries impact their workflow and, if necessary, exclude the small number of affected rows from their analysis.
When working with article text, we advise combining the title and subtitle fields into a single string. In some instances, one of the two fields may be missing or empty; merging them helps preserve as much article context as possible.
@@
Supporting code and example scripts to process the data can be found in the dataset folder [26]. Example code to join and decompress the embeddings files from the repository can be found in the README file.
We do not release the original web scraping code used to collect the data, as it involves platform-specific workarounds that may not apply at the current time and could place unnecessary strain on the servers of DerStandard. By sharing the processed dataset directly, we aim to reduce redundant scraping activity and minimize the load on the platform’s infrastructure.
Acknowledgements and Funding
We thank Elena Candellone, Darja Cvetković, Özgür Togay, David March, Timur Naushirvanov, Simon D. Lindner, and Elena G. de Lamo for testing the dataset during the Winter Workshop on Complex Systems 2025 and providing valuable feedback on its integrity and usability.

### v2

@@ Methods
As noted in the main text, we provide a separate file listing all comment identifiers associated with potential data inconsistencies. These include: (1) comments referencing article identifiers that are no longer available in the articles’ metadata (nor accessible on the DerStandard website), and (2) comments with timestamps preceding those of their associated articles. While these cases represent a small fraction of the data, they may affect specific analyses, such as estimating user lifespans. We recommend that users assess whether these entries impact their workflow and, if necessary, exclude the small number of affected rows from their analysis.
@@
Finally, researchers using this dataset should ensure that their analyses comply with applicable ethical standards for computational social science research, including avoiding attempts to re-identify individuals or infer sensitive personal attributes from the anonymized data.
Data Availability
The dataset described in this work is publicly available in the BSC Dataverse repository under the title “A Decade of News Forum Interactions: Threaded Conversations, Signed Votes, and Topical Tags” [36]. The dataset can be accessed via its DOI: https://doi.org/10.82201/RVQGUG. It is distributed under a CC-BY license and includes all data files, summary metadata, and example scripts required to access and process the dataset.
@@
Supporting code and example scripts to process the data can be found in the dataset folder Scripts [36], including two scripts demonstrating how to scrape forum discussions and votes from the platform at the time data was obtained. Example code to join and decompress the embedding files from the repository can be found in the README file.
Acknowledgements
We thank Elena Candellone, Darja Cvetković, Özgür Togay, David March, Timur Naushirvanov, Simon D. Lindner, and Elena G. de Lamo for testing the dataset during the Winter Workshop on Complex Systems 2025 and providing valuable feedback on its integrity and usability. We also thank the BSC Dataverse Data Stewards, Adrian Carrascosa Lopez and Felipe Leonardo Gomez Cortes, for their support and guidance in publishing the dataset in accordance with FAIR principles.
Funding